OpenStudy (anonymous):

Find the value of x & y hellppppp!!! x= 129 y 51 x 51 y 129 x66 y 17 x17 y 66

OpenStudy (anonymous):



OpenStudy (anonymous):

are those angles?

OpenStudy (anonymous):

2x-3=129 so 2x=132, x=66. 3y+129=180, 3y=51, y=51/3

OpenStudy (anonymous):

x=66 and y=17